MATEMATICA Anno Accademico 2005-2006 Scienze Geologiche M.Novaga, V.M. Tortorelli

————————————————————————————————————————–

I prova in itinere, 6 dicembre 2005 RISULTATI

————————————————————————————————————————–

ALEARDI Mattia  sufficiente
BACHINI Viola  valutabile per l’esonero
BELLONI Rossella  val.e.
BIONDI Giacomo  suf.
CAVALLARO Sharif  insufficiente
CHECCHI Giacomo  ins.
CIVITA Jacopo  ins.
FIGAIA Pietro  ins.
FRANCI Azzurra  ins.
FRISENDA Serena  ins.
GIANELLA Valentina  suf.
IACCARINO Salvatore  val.e.
LAURITI Elvira  ins.
MARCHETTI Samuele  discreto
NIRTA Antonio  val.e.
PANZIVO Egidio  ins.
PAOLINI Marco  val.e.+
PARDINI Dario  suf.
PASQUALONI Natascia  ins.
PELLEGRINI Andrea  val.e. +
PENNA Valentina  ins.
PIERETTI Carlo  suf.
RICCOMINI Matteo  val.e.
SANFIORENZO Angelo  ins.
SECCHIARI Arianna  val.e.
TAMBELLINI Simone  suf.
TENDOLI Simone  val.e.
VITI Federico  ins.